# Optimistic Query Routing in Clustering-based Approximate Maximum Inner Product Search

### v3

@@ 2. Routing by the Optimism Principle: our proposal
As we observed, NormalizedMean is an optimistic estimator, though its behavior is unpredictable. Our goal is to design an optimistic estimator that is statistically principled, thus well-behaved.
Before we begin, let us briefly comment on our terminology and notation. Throughout this section, we fix a unit query vector $q$; all discussions are in the context of $q$. We denote the $i$-th shard by $\mathcal{P}_{i}$, and write $S_{i}$ for the set of inner product scores between $q$ and points in $\mathcal{P}_{i}$: $S_{i}=\{\langle q,u\rangle:u\in\mathcal{P}_{i}\}$.
2.1. Formalizing the notion of optimism
@@
One can interpret the optimal $\theta_{i}$ as a probabilistic upper-bound on the maximum attainable value in $S_{i}$; that is, with some confidence, we can assert that no value in $S_{i}$ is greater than $\theta_{i}$.
Equipped with $\theta_{i}$’s, we route $q$ by sorting all shards by their estimated thresholds in descending order and selecting the top shards. Our router, dubbed Optimist, is defined as follow:
(4) $\tau(q)=\operatorname*{arg\,max\quad}^{(\ell)}_{i\in[C]}\theta_{i}.$
2.2. Understanding the routing behavior
Suppose for a moment that we have the solution to Problem 1, and let us expand on the expected behavior of the Optimist router. It is easy to see that, when $\delta\rightarrow 0$, the optimal solution approaches $\theta_{i}=\max(\langle q,\mu_{i}\rangle,\langle q,M_{i}\rangle)$, where $M_{i}$ is the median of the $i$-th partition. That represents the most conservative estimate of the maximum inner product between $q$ and points in $\mathcal{P}_{i}$.
As $\delta\rightarrow 1$, $\theta_{i}$ becomes larger, rendering Equation (4) an enthusiastically optimistic router. At the extreme, the optimal solution is the maximum inner product itself—the most optimistic we can get.
Clearly then, $\delta$ controls the amount of optimism one bestows onto the router. It is interesting to note that, when the data distribution is fully known, then $\delta=1$ is an appropriate choice: If we know the exact distribution of inner products, we can expect to be fully confident about the maximum inner product. On the other hand, when very little about the distribution is known (e.g., when all we know is the first moment of the distribution), then $\delta=0$ is a sensible choice. In effect, the value of $\delta$ is a statement on our knowledge of the underlying data distribution.
What is left to address is the solution to Problem 1, which is the topic of the remainder of this section. We first describe a general approach that uses as much or as little information as available about the data distribution. We then present a more practical approach that is the foundation of the rest of this work.
2.3. General solution
@@
Let us motivate our approach by considering a special case where $\mathcal{D}$ is $\mathcal{N}(\mu_{i},\Sigma_{i})$, a Gaussian with mean $\mu_{i}\in\operatorname*{\mathbb{R}}^{d}$ and covariance $\Sigma_{i}\in\operatorname*{\mathbb{R}}^{d\times d}$. In this case, $S_{i}$ follows a univariate Gaussian distribution with mean $\langle q,\mu_{i}\rangle$ and variance $q^{\top}\Sigma_{i}q$. In this setup, a solution to Problem 1 is simply:
(5) $\theta_{i}=\langle q,\mu_{i}\rangle+\sqrt{q^{\top}\Sigma_{i}q}\cdot\Phi_{\mathcal{N}(0,1)}^{-1}\left(\frac{1+\delta}{2}\right),$
which follows by expressing the cumulative distribution function (CDF) of $S_{i}$ in terms of the CDF of a unit Gaussian, denoted by $\Phi_{\mathcal{N}(0,1)}$.
Notice in the above example that, we first modeled the moments of $S_{i}$ using the moments of $\mathcal{D}$. We then approximated the CDF (equivalently, the quantile function) of the distribution of $S_{i}$ using its moments—in the Gaussian case, the approximation with the first two moments is, in fact, exact.
Our general solution follows that same logic. In the first step, we can obtain the first $r$ moments of the distribution of $S_{i}$, denoted by $m_{j}(S_{i})$ for $j\in[r]$, from the moments of $\mathcal{D}$, denoted by $m_{j}(\mathcal{D})$. In a subsequent step, we use $m_{j}(S_{i})$’s to approximate the CDF of $S_{i}$.
@@
In our second step, we wish to find a distribution $\tilde{S}_{i}$ such that $m_{j}(\tilde{S}_{i})\approx m_{j}(S_{i})$ for all $j\in[r]$. This can be done using the “method of moments,” a classic technique in statistics (Pearson, 1936) that offers guarantees (Kong and Valiant, 2017; Braverman et al., 2022) in terms of a distributional distance such as the Wasserstein-1 distance.
While using higher-order moments can lead to a better approximation, computing the $j$-th moment for $j>2$ can be highly expensive considering the dimensionality of datasets seen in practice, as the memory requirement to store the tensor $m_{j}(\mathcal{D})$ grows as $d^{j}$. We leave exploration and design of an efficient version of this two-step approach as future work.
2.4. Practical solution via concentration inequalities and sketching
@@
(6) $\theta_{i}=\langle q,\mu_{i}\rangle+\sqrt{\frac{1+\delta}{1-\delta}\;q^{\top}\Sigma_{i}q}.$
@@
$\Pr_{s\sim S_{i}}[s-\langle q,\mu_{i}\rangle\leq\epsilon]$ $\geq 1-\frac{q^{\top}\Sigma_{i}q}{q^{\top}\Sigma_{i}q+\epsilon^{2}}=\frac{1+\delta}{2}$
$\implies\epsilon^{2}=\frac{1+\delta}{1-\delta}\;q^{\top}\Sigma_{i}q.$
@@
We emphasize that the solution obtained by Lemma 1 is not necessarily optimal for Problem 1. Instead, it gives the best upper-bound on the optimal value of $\theta_{i}$ that can be obtained given limited information about the data distribution. As we see later, however, even this sub-optimal solution proves effective in practice.
While Equation (6) gives us an algorithm to approximate the threshold $\theta_{i}$, storing $\Sigma_{i}$ can be prohibitive in practice. That is because the size of the matrix grows as $d^{2}$ for each partition. Contrast that with the cost for other routers, such as Mean and NormalizedMean, which only store a single $d$-dimensional vector per partition (i.e., the mean vector). To remedy this inefficiency, we design a compact approximation of $\Sigma_{i}$ that can be plugged into Equation (6) to replace $\Sigma_{i}$. We describe that next to complete our algorithm.
2.4.1. Approximating the covariance matrix.
Since the procedure we describe is independently applied to each partition, we drop the subscript in $\Sigma_{i}$ and focus on a single partition.
We seek a matrix $\Sigma^{\ast}\in\operatorname*{\mathbb{R}}^{d\times d}$ that approximates the positive semi-definite (PSD) matrix $\Sigma$, by minimizing the following standard approximation error:
(7) $\text{err}(\Sigma,\Sigma^{\ast})\triangleq\sup_{\begin{subarray}{c}\|v\|_{2}=1\end{subarray}}|v^{\top}\Sigma v-v^{\top}\Sigma^{\ast}v|.$
As we seek a compact approximation, we constrain the solution space to rank-$t$ matrices: We require $\text{rank}(\Sigma^{\ast})=t$ for some $t\ll d$.
A standard solution to this problem is a result of the Eckhart-Young-Mirsky Theorem: Let $\Sigma=V\Lambda V^{\top}$ be the eigendecomposition of $\Sigma$. Then, $\Sigma^{\ast}=[V\sqrt{\Lambda}]_{t}[V\sqrt{\Lambda}]_{t}^{\top}$, where $[\cdot]_{t}$ is the operator that selects the first $t$ columns of its argument, optimally approximates $\Sigma$ under the stated constraints. We denote this solution by $\Sigma^{\textsc{lr}}_{t}$.
We could stop here and use $\Sigma^{\textsc{lr}}_{t}$ in lieu of $\Sigma$ in Equation (6). However, while that would be the optimal choice in the general case, for real-world datasets, we show that we can find a better approximation. In particular, we present the following:
@@
Rewrite $\Sigma$ as follows, with $D$ containing its diagonal and $R=\Sigma-D$:
@@
Let $Q\Lambda Q^{\top}$ be the eigendecomposition of $R_{\circ}$. We call the following the Masked Sketch of Rank $t$ of $\Sigma$:
(8) $\Sigma^{\textsc{ms}}_{t}=D+D^{\frac{1}{2}}[Q]_{t}[\Lambda]_{t}[Q]_{t}^{\top}D^{\frac{1}{2}}.$
The following result establishes that, under certain practical assumptions, $\text{err}(\Sigma,\Sigma^{\textsc{ms}}_{t})$ can be lower than $\text{err}(\Sigma,\Sigma^{\textsc{lr}}_{t})$.
Lemma 0.
Let $\Sigma\in\operatorname*{\mathbb{R}}^{d\times d}$ be a PSD matrix with diagonal $D$ for which $\min_{i\in[d]}{D_{ii}}/{\max_{i\in[d]}D_{ii}}\geq 1-\epsilon$ for some $\epsilon>0$. For every $11\leq t\leq d-$ 1 such that the $(t+1)$-th eigenvalue of $D^{-1/2}\Sigma D^{-1/2}$ is greater than $1$, we have that: $\text{err}(\Sigma,\Sigma^{\textsc{ms}}_{t})\leq\text{err}(\Sigma,\Sigma^{\textsc{lr}}_{t})\cdot\frac{1}{1-\epsilon}$.
We provide a proof for the above result in the next section and show that for real-world datasets used in this work, the assumptions in the lemma hold. But before we do so, we wish to recap our final solution.
@@
5:     $R_{\circ}=D^{-1/2}(\Sigma-D)D^{-1/2}$triangleright $D$ is the diagonal of $\Sigma$
@@
12:     $\tilde{q}\leftarrow q\circ\text{diag}(D^{1/2})$triangleright Element-wise product
@@
2.4.2. The final algorithm

### v4

@@ 2 Routing by the Optimism Principle: our proposal
As we observed, NormalizedMean is an optimistic estimator, though its behavior is unpredictable. Our goal is to design an optimistic estimator that is statistically principled, thus well-behaved. Throughout this section, all discussions are in the context of a fixed unit-vector $q$. We denote the $i$-th shard by $\mathcal{P}_{i}$, and write $S_{i}$ for the set of inner products of $q$ and points in $\mathcal{P}_{i}$: $S_{i}=\{\langle q,u\rangle:u\in\mathcal{P}_{i}\}$.
2.1 Formalizing the notion of optimism
@@
The optimal $\theta_{i}$ is a probabilistic upper-bound on the maximum attainable value in $S_{i}$; that is, with some confidence, we can assert that no value in $S_{i}$ is greater than $\theta_{i}$. We then route $q$ by sorting $\mathcal{P}_{i}$’s by their $\theta_{i}$ in descending order and select the top shards. Our router, Optimist, is defined as follow:
$\tau(q)=\operatorname*{arg\,max\quad}^{(\ell)}_{i\in[C]}\theta_{i}.$ (4)
Suppose that we have the solution to Problem 1. As $\delta\rightarrow 0$, the optimal solution approaches $\theta_{i}=\max(\langle q,\mu_{i}\rangle,\langle q,M_{i}\rangle)$, where $M_{i}$ is the median of $\mathcal{P}_{i}$. That is a conservative estimate of the maximum inner product. As $\delta\rightarrow 1$, $\theta_{i}$ becomes larger, rendering Equation (4) optimistic. At the extreme, the optimal $\theta_{i}$ is the maximum inner product itself—the most optimistic we can get.
Clearly $\delta$ controls the degree of optimism. When the distribution is fully known, $\delta=1$ is an appropriate choice: If we know the distribution of inner products, we can expect to be fully confident about the maximum inner product. But when very little about the distribution is known, then $\delta=0$ is a sensible choice. In effect, the value of $\delta$ reflects our knowledge of the data distribution.
What is left to address is the solution to Problem 1, which is the topic of the remainder of this section. We first describe a general approach that uses as much information as available about the data distribution. We then present a practical approach that is the foundation of the rest of this work.
2.2 General solution
@@
We motivate our approach by considering a special case where $\mathcal{D}$ is $\mathcal{N}(\mu_{i},\Sigma_{i})$, a Gaussian with mean $\mu_{i}\in\operatorname*{\mathbb{R}}^{d}$ and covariance $\Sigma_{i}\in\operatorname*{\mathbb{R}}^{d\times d}$. In this case, $S_{i}$ follows a univariate Gaussian distribution with mean $\langle q,\mu_{i}\rangle$ and variance $q^{\top}\Sigma_{i}q$, so that the solution to Problem 1 is simply:
$\theta_{i}=\langle q,\mu_{i}\rangle+\sqrt{q^{\top}\Sigma_{i}q}\cdot\Phi_{\mathcal{N}(0,1)}^{-1}\left(\frac{1+\delta}{2}\right),$ (5)
which follows by writing the cumulative distribution function (CDF) of $S_{i}$ in terms of the CDF of a unit Gaussian, denoted $\Phi_{\mathcal{N}(0,1)}$. Notice that, we first modeled the moments of $S_{i}$ using the moments of $\mathcal{D}$, then approximated the CDF (equivalently, the quantile function) of the distribution of $S_{i}$ from its moments—in this special case, the approximation with the first two moments is, in fact, exact.
Our general solution follows that same logic. In the first step, we can obtain the first $r$ moments of the distribution of $S_{i}$, denoted $m_{j}(S_{i})$ for $j\in[r]$, from the moments of $\mathcal{D}$, denoted $m_{j}(\mathcal{D})$. In a subsequent step, we use $m_{j}(S_{i})$’s to approximate the CDF of $S_{i}$.
@@
In a second step, we find a distribution $\tilde{S}_{i}$ such that $m_{j}(\tilde{S}_{i})\approx m_{j}(S_{i})$ for all $j\in[r]$. This can be done using the “method of moments” Pearson (1936) which offers guarantees (Kong and Valiant, 2017; Braverman et al., 2022) in terms of a distributional distance such as the Wasserstein-1 distance.
While using higher-order moments leads to a better approximation, computing $m_{j}(\cdot)$ for $j>2$ can be prohibitive considering the dimensionality of real datasets, as the space requirement to store $m_{j}(\mathcal{D})$ grows as $d^{j}$. We leave an exploration of an efficient version of this two-step approach as future work.
2.3 Practical solution via concentration inequalities and sketching
@@
$\theta_{i}=\langle q,\mu_{i}\rangle+\sqrt{\frac{1+\delta}{1-\delta}\;q^{\top}\Sigma_{i}q}.$ (6)
@@
$\Pr_{s\sim S_{i}}[s-\langle q,\mu_{i}\rangle\leq\epsilon]$ $\geq 1-\frac{q^{\top}\Sigma_{i}q}{q^{\top}\Sigma_{i}q+\epsilon^{2}}=\frac{1+\delta}{2}\implies\epsilon^{2}=\frac{1+\delta}{1-\delta}\;q^{\top}\Sigma_{i}q.$
@@
We emphasize that Lemma 1 gives a loose upper-bound on the optimal value of $\theta_{i}$. Because we make no assumptions about the data distribution, the gap between the optimal and estimated $\theta_{i}$ cannot be expressed in closed form. Had we assumed $S_{i}$ is sub-Gaussian, it would be easy to bound the gap between $\theta_{i}$ obtained from Lemma 1 and the one obtained from sub-Gaussian concentration bounds. Nonetheless, we find that in practice even this sub-optimal solution proves effective.
There is still one challenge with Equation (6): $\Sigma_{i}$ can be too large to store as the matrix grows as $d^{2}$ for each partition. Contrast that with the cost of Mean and NormalizedMean which only need one $d$-dimensional vector per partition. To remedy this, we design a compact approximation of $\Sigma_{i}$ that can be plugged into Equation (6) to replace $\Sigma_{i}$. We describe that next to complete our algorithm.
2.3.1 Approximating the covariance matrix
Since the procedure we describe is independently applied to each partition, we drop the subscript in $\Sigma_{i}$ and focus on a single partition. We seek a matrix $\Sigma^{\ast}\in\operatorname*{\mathbb{R}}^{d\times d}$ that approximates the positive semi-definite (PSD) matrix $\Sigma$, by minimizing the following standard approximation error:
$\text{err}(\Sigma,\Sigma^{\ast})\triangleq\sup_{\begin{subarray}{c}\|v\|_{2}=1\end{subarray}}|v^{\top}\Sigma v-v^{\top}\Sigma^{\ast}v|.$ (7)
As we seek a compact $\Sigma^{\ast}$, we constrain the solution space to rank-$t$ matrices: $\text{rank}(\Sigma^{\ast})=t$, $t\ll d$.
A standard solution is a result of the Eckhart-Young-Mirsky Theorem: Let $\Sigma=V\Lambda V^{\top}$ be the eigendecomposition of $\Sigma$. Then, $\Sigma^{\ast}=[V\sqrt{\Lambda}]_{t}[V\sqrt{\Lambda}]_{t}^{\top}$, where $\cdot]_{t}$ selects the first $t$ columns of its argument, optimally approximates $\Sigma$ under the stated constraints. We denote this solution by $\Sigma^{\textsc{lr}}_{t}$.
We could stop here and use $\Sigma^{\textsc{lr}}_{t}$ in lieu of $\Sigma$ in Equation (6). However, while that would be the optimal choice in the general case, for real-world datasets, we show that we can find a better approximation.
@@ 2.3.1 Approximating the covariance matrix
Rewrite $\Sigma$ as follows, with $D$ its diagonal and $R=\Sigma-D$:
@@
Let $Q\Lambda Q^{\top}$ be $R_{\circ}$’s eigendecomposition. We call the following the Masked Sketch of Rank $t$ of $\Sigma$:
$\Sigma^{\textsc{ms}}_{t}=D+D^{\frac{1}{2}}[Q]_{t}[\Lambda]_{t}[Q]_{t}^{\top}D^{\frac{1}{2}}.$ (8)
@@
5:  $R_{\circ}=D^{-1/2}(\Sigma-D)D^{-1/2}$ $\triangleright$ $D$ is the diagonal of $\Sigma$
@@
12:  $\tilde{q}\leftarrow q\circ\text{diag}(D^{1/2})$ $\triangleright$ Element-wise product
@@
We next establish that, under certain assumptions, $\text{err}(\Sigma,\Sigma^{\textsc{ms}}_{t})$ is lower than $\text{err}(\Sigma,\Sigma^{\textsc{lr}}_{t})$.
Lemma 2.
Let $\Sigma\in\operatorname*{\mathbb{R}}^{d\times d}$ be a PSD matrix with diagonal $D$ for which $\min_{i\in[d]}{D_{ii}}/{\max_{i\in[d]}D_{ii}}\geq 1-\epsilon$ for some $\epsilon>0$. For every $1\leq t\leq d-1$ such that the $(t+1)$-th eigenvalue of $D^{-1/2}\Sigma D^{-1/2}$ is greater than $1$, we have that: $\text{err}(\Sigma,\Sigma^{\textsc{ms}}_{t})\leq\text{err}(\Sigma,\Sigma^{\textsc{lr}}_{t})/(1-\epsilon)$.
We give the proof in Appendix C and show that for datasets in this work the assumptions hold.
2.3.2 The final algorithm